OpenStudy (anonymous):

Find critical points from dy/dx = yln(y+2). I know that we have to set: yln(y+2) = 0, but don't know what to do to get started after that....

OpenStudy (primeralph):

What's the original question?

OpenStudy (anonymous):

find the critical points and phase portrait of the given autonomous first-order differential equation. Classify each critical point as asymptotically stable, unstable, or semi-stable.

OpenStudy (anonymous):

I just can't get past the first part where it say s critical points....

OpenStudy (anonymous):

the critical points would be the y values that make the equation y(ln(y+2)=0 or if the y values make the equation impossible

OpenStudy (anonymous):

I can't figure that out with ln..

OpenStudy (anonymous):

it should be y=0 and y=-2

OpenStudy (anonymous):

ln(0) does not exist therefore it is a critical point

OpenStudy (ybarrap):

yln(y+2) = 0 when y = 0 or ln(y+2) = 0 or y = -1

OpenStudy (anonymous):

why is y=-1 a critical point

OpenStudy (anonymous):

oh nvm

OpenStudy (ybarrap):

to find if stable or unstable, take deriviative of the yln(y+2), evaluated at each critical point and check if > 0 or < 0. If < 0 then stable, if > 0 unstable

OpenStudy (anonymous):

thanks.